Monochromatic light of wavelength 592 nm from a distant source passes through a slit that is 0.0290 mm wide. In the resulting diffraction pattern, the intensity at the center of the central maximum (θ= 0°) is 4.00 x 10-5 W/m2. What is the intensity at a point on the screen that corresponds to θ = 1.20°?

1
Expert's answer
2022-01-12T08:31:38-0500
"I=I_0\bigg(\dfrac{\sin\theta}{\theta}\bigg)^2=1.22·10^{-8}\text{ W/m}^2."